# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

### v2

@@ 2.3.2 Markov Chain Monte Carlo Sampling
Figure 6: Example of the MCMM application on the inverse inference of cardiac activation properties from ECG. Here, MCMC has been used for the sampling of parameter sets of the population [103]. Image adapted from Camps et al. [103] with permission.
Current probabilistic inference approaches are generally based on Markov Chain Monte Carlo (MCMC) sampling methods to handle complex probabilistic models and uncertainty associated with ECG data [104]. For instance, Camps et al. [103] presented a sequential Monte Carlo approximate Bayesian computation-based model for the inference of ventricular activation properties. Rahimi et al. [105] applied a hierarchical Bayesian inference to estimate the statistical parameters in the posterior pdf, which can be calculated via MCMM sampling methods. However, direct MCMC sampling of the pdf of the parameters can be impractical, as it involves repeated evaluations of the posterior pdf which consists of computationally expensive simulation model. To accelerate sampling, various hybrid sampling techniques have been developed using information about the target pdf such as gradient and Hessian matrix [106], which however are challenging to extract when dealing with complex simulation models. Alternatively, one could develop computationally efficient surrogate models of the expensive simulation process, making it significantly faster to sample the associated pdfs [107, 108]. Despite improved efficiency, direct sampling of surrogate-based posterior pdf may yield limited accuracy, given the challenge of creating accurate approximations for complex nonlinear simulation models. Instead of completely replacing the sampling of the exact posterior pdf, Dhamala et al. [83] proposed a two-stage model to integrate Gaussian process surrogate modelling of the posterior pdf in accelerating sampling. Also, one could directly develop a surrogate model for the posterior pdf of simulation model parameters, eliminating the necessity for additional MCMC sampling of the computationally intensive original pdf [104].
@@ 2.3.3 Variational Inference
Variational inference (VI) approximates the posterior distribution with a simpler variational distribution by solving an optimization problem. This optimization process is often computationally more efficient than MCMC sampling as it does not require repeated sampling from the posterior distribution. Therefore, the application of VI has been widely adopted in solving the ECG inverse problem [74]. For instance, Xu et al. [99] introduced a Bayesian inference framework for robust transmural electrophysiological imaging. Ghimire et al. [74] proposed a Bayesian model for the estimation of the prior model error while reconstructing TMP under the constraint of a 3D EP simulation model. These conventional VI methods are generally optimized using techniques such as gradient descent and expectation-maximization (EM) algorithms.
Recently, deep learning-based probabilistic methods have served as an emerging alternative to conventional methods for modelling complex dynamics of cardiac electrical activity [109, 110, 111]. They can leverage deep neural networks to approximate the posterior distribution of the model parameters or latent variables, providing faster and more accurate approximations. For example, Ghimire et al. [109] proposed a deep generative model to reconstruct cardiac transmembrane potential from ECG data. Bacoyannis et al. [112] employed a variational auto-encoder (VAE) model to reconstruct the activation pattern of the myocardium with various local wall thicknesses as thin walls indicated infarct areas. Bacoyannis et al. [113] employed conditional VAE to generate activation maps from combined BSP and conductivity maps. Li et al. [114] designed a VAE-based deep computational model for the inverse inference of ventricular activation properties, i.e., root node (RN) position and CVs. Dhamala et al. [115] employed a graph convolutional VAE to allow generative modelling of non-Euclidean data for cardiac model personalization. Dhamala et al. [110] proposed a generative VAE for parameter estimation of a personalized cardiac model. Yang et al. [111] proposed Bayesian PINN to solve linear or nonlinear PDEs with noisy data for both forward and inverse problems. They employed the Hamiltonian Monte Carlo or the variational inference for the estimation of the posterior distributions.
Figure 7: Example of the conditional-VAE model for the inverse inference of myocardial infarction (MI) [116]. Image adapted from Li et al. [116] with permission.
@@ 2.3.4 Surrogate Model for Bayesian Optimization
To accelerate the inverse inference of parameters, surrogate models, especially statistical emulation with Gaussian processes (GP), have been employed in cardiovascular fluid dynamics [117], the pulmonary circulatory system [118], ventricular mechanics [119], and LV cardiac dynamics [117]. It involves approximating the computationally intensive mathematical model (the simulator) with a computationally efficient statistical surrogate model (the emulator) [117]. This is achieved through a combination of extensive parallelization and GP regression, allowing for optimal utilization of computational resources prior to the patient arrival at the clinic. Here, GP regression is a non-parametric Bayesian approach used for regression tasks where the model learns from the observed data to make predictions. It does not assume a specific parametric form for the underlying function but instead models the function as a distribution over functions, which is determined by the observed data points. Therefore, when new data are accessible, the proxy objective function can be minimized at a low computational cost. This eliminates the need for additional computationally intensive simulations from the original mathematical model and thus accelerates the inference. Recently, the GP surrogate model started to be used in the ECG inverse inference by Pezzuto et al. [120], where the GP surrogate model with Bayesian optimization has been used for the inverse inference of RNs. These strategies offer a mid-way solution between pure sampling methods using mechanistic models and machine learning-based solutions as the variational inference. They retain the flexibility and interpretability of pure sampling methods while replacing the mechanistic model with a faster emulator, which can be machine learning-based.
3 Uncertainty Quantification for the ECG Inverse Problem
Solving the ECG inverse problem relies on many subject-specific measurements and modelling assumptions, which introduce various sources of uncertainty in the predictions. Uncertainty quantification aims to analyze how models respond to inherent variability in input parameters. ECG inverse problems are often ill-posed and lack uniqueness in their solutions. However, clinicians typically require a single prediction for practical decision-making purposes in patient care scenarios. Therefore, it is essential to quantify and mitigate the effects of different types of uncertainties to ensure the robustness and clinical applicability of ECG inverse solutions.
3.1 Measurement Acquisition Uncertainty
ECG inverse inference requires subject-specific EP measurements, which can be easily distorted by noise artifacts, such as motion artifacts. EP data typically exhibit sparsity, with measured values associated with a higher degree of uncertainty [121]. For example, noise, a ubiquitous challenge in real-world measurements, can significantly distort the recorded ECG signals, introducing uncertainty into the source localization. To reduce the effects of measurement noise on the inverse problem, Aydin et al. [101] used a Kalman filter-based approach. Dogrusoz et al. [122] employed Bayesian estimation for solving the inverse ECG problem by applying various prefiltering methods to reduce the effects of noise. Similarly, Neumann et al. [123] utilized stochastic parameter estimation, aggregating probabilities estimated under various noise levels to derive a robust parameter estimate without explicitly determining the level of noise in the measurement. Moreover, there is ambiguity in matching the measurement points in the BSPM recording with the computational mesh derived from imaging data due to the non-synchronized BSPM-imaging acquisition [107]. The heart position is also susceptible to various sources of movement artifacts, including respiration and cardiac contraction, introducing geometric errors into the inverse problem of ECG solutions [124]. Therefore, the geometric error in the anatomical modelling (forward model) should also be considered, as it will in turn cause estimation errors in the ECG inverse inference.
3.2 Anatomical Modeling Uncertainty
The anatomical modelling is normally started with image segmentation, and then numerical solution techniques will be used to solve the forward ECG problem [58]. Therefore, the precision of the anatomical model employed, hinging on segmentation accuracy [125, 126], inter-subject anatomical variations [127], structural remodelling of the myocardium [116], and mesh resolution [128], profoundly influences the reliability of inverse solutions. Incorporating a prior model can mitigate uncertainties, and a combination of prior models may be preferred for estimating complex source structures [105]. Also, one can disentangle the inter-subject anatomical variation to mitigate its effect on the ECG inverse problem [127]. Furthermore, biventricular pacing and scar size, often encountered in clinical scenarios, demand meticulous modelling for a faithful representation [129]. Besides the heart, the torso geometry accuracy can also subtly influence the inverse solution, as the EP measurement is generally performed on the body surface [130]. In most studies, each torso surface node often corresponds to an electrode, and thus insufficient electrodes may lead to non-representative mesh nodes, requiring interpolation techniques for solving the inverse problem [131]. Dogrusoz et al. [131] investigated the effects of interpolation on the inverse problem of ECG. There also exist many studies exploring the effects of cardiac position inside the torso on the inverse inference [132, 133]. To address uncertainties, advanced regularization techniques that encompass prior knowledge, noise reduction strategies, and advanced data assimilation methods show promise [58, 134]. As the field advances, careful consideration of these effects becomes paramount for robust and clinically meaningful results in solving the ECG inverse problem.
Figure 8: The cardiac digital twin generation workflow combining cardiac imaging and ECG data, with ECG inverse inference for personalization. Image modified from Li et al. [116] with permission.
3.3 Computational Modeling Assumption Uncertainties
Model assumptions are necessary simplifications or idealizations of reality that are made to facilitate mathematical or computational modelling. These assumptions may include assumptions about tissue properties, boundary conditions, geometry, and physiological processes. In ECG inverse inference, assumptions are often made about the homogeneity of tissue properties, the behavior of electrical signals within the heart, and the relationship between measured ECG signals and underlying electrical activity [126, 135, 136]. While model assumptions are essential for simplifying complex systems, they can also introduce uncertainties and errors into the modelling process. For example, assuming uniform tissue properties throughout the heart and torso may lead to inaccuracies in the predicted electrical activity, especially in regions with significant heterogeneity [130, 53]. Similarly, assuming specific boundary conditions or physiological behaviors may not fully capture the complexity of real-world scenarios, leading to discrepancies between model predictions and observed data. ECG inverse inference also depends heavily on the assumptions on the initial distribution of heart electrical potential [135]. Determining the initial potential distribution at the onset of the cardiac cycle is particularly challenging due to the difficulty in accurately pinpointing the anatomical site of the earliest excitation in the heart, often associated with the intricate Purkinje networks [95]. Therefore, sophisticated modelling strategies are required to adequately account for the influence of these networks on ECG inverse inference [137, 138]. Once uncertainties are quantified, strategies can be employed to mitigate their impact on model predictions. The process of balancing model assumptions and uncertainty is often iterative, requiring continuous refinement and validation of models against experimental or clinical data. This iterative refinement process helps researchers to build more robust and reliable models over time.
@@ 4.1 Dataset
The solution to the ECG inverse inference problem is typically implemented within a cardiac digital twin platform, as shown in Fig. 8. Therefore, for evaluation purposes, access to both anatomical data from cardiac imaging and electrical activity recording data from ECG or BSP mapping is necessary. While there are many public cardiac imaging data [139, 140, 141, 142, 143, 144] and ECG recordings [145, 146, 147, 148], the evaluation of ECG inverse inference methods requires paired imaging and ECG/ BSP mapping data from the same patient, which is less publicly available. UK Biobank (UKB) [149] and China Kadoorie Biobank (CKB) [150, 151] provide paired imaging and ECG dataset. MyoFit46 includes 500 participants of approximately 75 years+ to undertake high-density surface ECGI and cardiac MRI [152]. Strocchi et al. [153] provided a publicly available virtual cohort of twenty-four four-chamber hearts built from heart failure patients. Experimental Data and Geometric Analysis Repository (EDGAR) is another public dataset for the application and validation of ECGI techniques [154]. It includes paired BSPM and endocardial/ epicardial recordings from both animals (dog, pig, etc) and human and torso tank datasets. Additionally, information regarding the cardiac position within the torso is essential, often requiring torso image data. Several studies have attempted to reconstruct the torso directly from standard cardiac imaging for the localization of electrodes, which was challenging due to partial torso shape information [155, 156]. There are also many validation studies conducted in ex vivo torso tanks [157, 158, 159], in vivo large animal models [160, 161, 93], or in patients [162, 25]. Fig. 9 and Fig. 10 provide the two most commonly used ex vivo torso tanks.
Figure 10: Illustration of Bordeaux Torso Tank, with a pig heart perfused in Langendorff mode with a 108-electrode sock placed over the ventricles. Image adapted from Bear et al. [158] with permission.
Figure 11: Illustration of different cardiac source models in ECG.
@@ 4.2 Evaluation
The validation of the inverse problem involves a diverse set of cardiac source models, forward EP model formulations, and inverse methods, as presented in Table 3. For simplicity, we organized the validation according to the estimated cardiac source models, including surface potential models, activation/ recovery models, and transmembrane voltage models [52]. Fig. 11 provides three representative examples corresponding to the three source models. Except for a few studies, all the others employed BSP as the input representation/ measurement for cardiac electrical activity assessment. One can see that most studies employed computer models for in silico evaluation [66, 163] and/ or performed evaluation on the basis of torso-tank experiments with isolated animal hearts, such as rabbit [164], canine [77, 73, 92], and swine [165, 166, 167]. There are also few studies evaluated using in-vivo human subjects, which however usually employed non-simultaneous invasive recordings, such as EGMs [168, 169] or bipolar voltages maps [81, 74]. The comparison with the published invasive measurements in animals or humans has also been utilized as an alternative evaluation [24].
Table 3: Summary of the ECG inference works. EP: electrophysiology; AS: activation sequence; EGM: electrograms; HF: heart failure; HS: healthy subject; EpiP/ EndoP: epicardial/ endocardial potential; RN: root node; VCG: vectorcardiographic; AP: Aliev-Panfilov; LSTM: long short-term memory; FCN: fully connected neural network; TV: total-variation.
@@
Source & Publish year Data Input Target EP model Inverse inference method
Ramanathan et al. (2004) [24] 1 HS; 3 patients BSP EpiP, EGM, isochrones boundary element Tikhonov regularization + generalized minimal residual algorithm
@@
Shou et al. (2008) [60] in silico BSP EpiP finite element truncated total least squares
Wang et al. (2009) [63] in silico BSP EpiP finite element Tikhonov regularization
Jiang et al. (2009) [170] in silico BSP EpiP boundary element subspace preconditioned LSQR
Sarikaya et al. (2010) [59] 1 canine BSP EpiP N/A genetic algorithm with conventional regularization
Wang et al. (2010) [171] 1 canine BSP EpiP finite element variational-form-based regularizers
Wang et al. (2011) [172] 1 canine BSP EpiP finite element variational regularizers
Cluitmans et al. (2013) [168] 1 patient BSP EpiP N/A realistic data and VCG based optimization
Zemzemi et al. (2013) [173] in silico BSP EpiP bidomain Steklov-Poincare formulation + gradient descent method
Zemzemi et al. (2014) [174] in silico BSP EpiP bidomain iterative Kozlov-Maz’ya-Fomin
@@
Erem et al. (2014) [82] in silico; 1 HS BSP EpiP, EndoP N/A temporal spline model
Yao et al. (2017) [175] 4 MI patients BSP EGM, MI N/A spatiotemporal regularization
Cluitmans et al. (2017) [176] 3 canine BSP EpiP AP physiology-based regularization
Cluitmans et al. (2018) [73] 3 canine BSP EpiP, ATM, beat origin N/A multitask elastic-net-based method
Kalinin et al. (2019) [163] in silico BSP EpiP, EndoP bidomain single layer approach
@@
Karoui et al. (2019) [88] in silico BSP EpiP monodomain spatial adaptation of time delay neural network
Bear et al. (2020) [65] 4 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
Rababah et al. (2021) [67] 6 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
Ozkoc et al. (2021) [177] 1 canine BSP EpiP boundary element Bayesian MAP
Chen et al. (2022) [167] 5 swines ECG EpiP, ATM, RN N/A FCN, CNN, LSTM model
Tenderini et al. (2022) [85] in silico ECG EpiP AP PDE-aware deep learning model
Jiang et al. (2022) [92] 1 canine BSP EpiP AP physics embedded graph CNN
Xie et al. (2023) [178] in silico BSP EpiP AP physics-constrained network
@@
Jiang et al. (2024) [179] in silico; 3 patients BSP EpiP, EndoP reaction-Eikonal model neural state-space modeling
Messnarz et al. (2004) [68] in silico BSP TMP bidomain spatiotemporal regularization
Wang et al. (2009) [79] in silico BSP TMP bidomain physiological-model-constrained statistical approach
Wang et al. (2011) [180] in silico BSP TMP, MI bidomain physiologically-based constraints
Potyagaylo et al. (2012) [181] in silico BSP TMP N/A iterative Gauss-Newton optimization
Jiang et al. (2013) [182] in silico BSP TMP bidomain Support vector regression
Lopez et al. (2014) [183] in silico BSP TMP monodomain genetic algorithm with Tikhonov regularization
Xu et al. (2014) [81] in silico; 6 patients BSP MI N/A variational Bayesian with TV prior
@@
Zettinig et al. (2014) [38] in silico ECG Myo electrical diffusivity LB polynomial regression
Ghimire et al. (2019) [74] in silico; 2 patients ECG TMP finite element Bayesian model
Kara et al. (2019) [64] in silico BSP TMP, MI bidomain Tikhonov regularization
Cheng et al. (2021) [184] in silico BSP TMP, MI bidomain shrinkage-thresholding network
Mu et al. (2021) [185] in silico BSP TMP N/A shrinkage-thresholding + graph CNN
Zhao et al. (2022) [186] in silico BSP TMP bidomain data-driven Kalman filtering network
Li et al. (2024) [116] in silico ECG MI Eikonal multi-modal VAE
Lian et al. (2024) [187] in silico; 4 patients ECG MI bidomain frequency-enhanced model
Zhang et al. (2005) [164] 4 rabbits BSP ATM N/A N/A
Calderero et al. (2005) [77] 1 canine BSP activation wavefront potential-based level sets
Alawad et al. (2019) [188] in silico; 3 patients ECG pacing sites N/A L1-norm regularization
Schuler et al. (2021) [66] in silico BSP ATM bidomain Tikhonov regularization
Pezzuto et al. (2021) [169] 11 HF patients ECG ATM, RN Eikonal propagation model
Meister et al. (2021) [165] 16 swine ECG ATM N/A graph convolutional regression
Camps et al. (2021) [103] in silico ECG CV, RN N/A Monte Carlo Bayesian inf
Li et al. (2022) [189] in silico ECG CV, RN Eikonal conditional VAE

### v4

@@ 2.3.3 Markov Chain Monte Carlo Sampling
Probabilistic estimation of EP parameters is crucial for model personalization and uncertainty quantification. Many probabilistic inference methods use Markov Chain Monte Carlo (MCMC) sampling to handle complex models and uncertainties associated with ECG data [136]. For instance, Camps et al. [1] presented a sequential Monte Carlo approximate Bayesian computation-based model for the inference of ventricular activation properties, as shown in Fig. 4. Rahimi et al. [137] applied a hierarchical Bayesian inference to estimate the statistical parameters in the posterior pdf, which can be calculated via MCMM sampling methods. However, direct MCMC sampling of the pdf of the parameters can be impractical, as it involves repeated evaluations of the posterior pdf involving computationally expensive simulation processes. To accelerate sampling, various hybrid sampling techniques have been developed using information about the target pdf such as gradient and Hessian matrix [138], which however are challenging to extract when dealing with complex simulation models. Alternatively, one could develop computationally efficient surrogate models of the expensive simulation process, making it significantly faster to sample the associated pdfs [139, 140]. Despite improved efficiency, direct sampling of surrogate-based posterior pdf may yield limited accuracy, given the challenge of creating accurate approximations for complex nonlinear simulation models. Instead of completely replacing the sampling of the exact posterior pdf, Dhamala et al. [141] proposed a two-stage model to integrate Gaussian process surrogate modeling of the posterior pdf in accelerating sampling. Also, one could directly develop a surrogate model for the posterior pdf of simulation model parameters, eliminating the necessity for additional MCMC sampling of the original computationally-intensive pdf [136].
@@ 2.3.4 Variational Inference
For ECGI, unlike conventional regularization techniques, variational inference (VI) provides a probabilistic framework that can quantify uncertainty in the reconstructed potentials. Specifically, it can incorporate prior knowledge about the spatial and temporal dynamics of cardiac activity, improving the robustness of the solution against noisy data. For example, Ghimire et al. [142] introduced a Bayesian framework to jointly infer the posterior distribution of TMP and the errors in the prior EP model from ECG data under the constraint of a 3D EP simulation model. Xu et al. [143] introduced a Bayesian inference framework for robust transmural electrophysiological imaging. Recently, deep learning-based probabilistic methods have served as an emerging alternative to conventional methods for modeling complex dynamics of cardiac electrical activity [144, 145]. They can leverage deep neural networks to approximate the posterior distribution of the model parameters or latent variables, providing faster and more accurate approximations. Specifically, variational auto-encoder (VAE), as a extension of traditional VI, is quite useful to handle complex, high-dimensional data during the inverse inference. For example, Bacoyannis et al. [102] employed a VAE model to reconstruct the activation pattern of the myocardium with various local wall thicknesses as thin walls indicated infarct areas. Bacoyannis et al. [146] employed conditional VAE to generate activation maps from combined BSP and conductivity maps.
For EP parameter estimation, VI approximates the posterior distribution of EP parameters with a simpler variational distribution by solving an optimization problem. Similar to ECGI, VAE has also been widely used for EP parameter estimation, such as the inverse inference of ventricular activation properties, i.e., EASs and CVs [147]. Graph convolutional VAE can further enhance this estimation by effectively handling non-Euclidean data structures, i.e., cardiac meshes, allowing for more accurate modeling of spatial relationships within the heart [148]. Specifically, Dhamala et al. employed graph convolutional VAE to embed the high-dimensional optimization into low-dimensional latence space for the estimation of tissue excitability [149]. Recently, Yang et al. [145] proposed Bayesian PINN to solve linear or nonlinear PDEs with noisy data for both forward and inverse problems. They employed the Hamiltonian Monte Carlo or the variational inference for the estimation of the posterior distributions. This approach could be effectively adapted to tackle the ECG inverse problem in the future.
@@ 2.3.5 Surrogate Model for Bayesian Optimization
To accelerate the inverse inference of parameters, surrogate models, especially statistical emulation with Gaussian processes (GP), have been employed in cardiovascular fluid dynamics [150], the pulmonary circulatory system [151], ventricular mechanics [152], and LV dynamics [150]. It involves approximating the computationally intensive mathematical model (the simulator) with a computationally efficient statistical surrogate model (the emulator) [150]. This could be achieved through a combination of extensive parallelization and GP regression, allowing for optimal utilization of computational resources prior to the patient arrival at the clinic. GP regression is a non-parametric Bayesian approach used for regression tasks where the model learns from the observed data to make predictions. It does not assume a specific parametric form for the underlying function but instead models the function as a distribution over functions, which is determined by the observed data points. Therefore, when new data are accessible, the proxy objective function can be minimized at a low computational cost. This eliminates the need for additional computationally intensive simulations from the original mathematical model and thus accelerates the inference. Recently, the GP surrogate model started to be used in the ECG inverse inference by Pezzuto et al. [153], where the GP surrogate model with Bayesian optimization has been used for the estimation of EASs. Earlier work by Konukoglu et al. [139] demonstrated Bayesian estimation using polynomial chaos for EASs and conductivity estimation. These strategies offer a mid-way solution between pure sampling methods using EP models and machine learning-based solutions as the variational inference. They retain the flexibility and interpretability of pure sampling methods while replacing the model with a faster emulator, which can be machine learning-based.
@@ 3 Dataset and Evaluation
Figure 6: Illustration of in-vivo evaluation based on the rabbit undergoing the ventricular pacing from various locations. Illustrations designed referring to Han et al. [95].
@@ 3.1 Dataset
To evaluate the solution of ECG inverse problem, both anatomical data from cardiac imaging and electrical activity recording data from ECG or BSP are necessary, as shown in Fig. 1. While there are many public cardiac imaging data [154, 155, 156, 157] and ECG recordings [158, 159, 160], the evaluation of ECG inverse inference methods requires paired imaging and signal data from the same patient, which is less publicly available. UK Biobank [161] and China Kadoorie Biobank [162, 163] provide paired imaging and ECG dataset. MyoFit46 includes 500 participants of approximately 75 years+ to undertake high-density surface ECGI and cardiac MRI [164]. Strocchi et al. [165] provided a publicly available virtual cohort of twenty-four four-chamber hearts built from heart failure patients. Experimental Data and Geometric Analysis Repository (EDGAR) is another public dataset for the application and validation of ECGI techniques [166]. It includes paired BSPM and endocardial/ epicardial recordings from both animals and torso tank datasets. Accurate cardiac positioning within the torso is also crucial and often requires torso imaging data for precise localization. Several studies have attempted to reconstruct the torso directly from standard cardiac imaging for the localization of electrodes, which was challenging due to partial torso shape information [167, 168, 169]. There are also several validation studies conducted in ex vivo torso tanks [36, 170, 171], in vivo large animal models (rabbit, canine, swine, etc) [172, 173, 84, 95, 174], or in human [175, 26]. Fig. 5 provides the one of the most commonly used ex-vivo torso tanks. Fig. 6 presents an in-vivo rabbit model for the evaluation of non-invasive 3D activation sequence reconstruction.
@@ 3.2 Evaluation
Table 3: Summary of selected ECG inverse inference works for ECGI. EGM: electrograms; OVA: origin of ventricular activation; VT: ventricular tachycardia; MI: myocardial infarction; AF: atrial fibrillation; PVC: premature ventricular contraction; EpiP/ EndoP: epicardial/ endocardial potential.
@@
Source & Publish year Data Signal # Electrodes Target Inverse inference method
@@
Deterministic estimation Shou et al. (2008) [56] in silico BSP N/A EpiP truncated total least squares
Wang et al. (2009) [59] in silico BSP N/A EpiP Tikhonov regularization
Wang et al. (2013) [176] in silico BSP 400 Myo ischemia, TMP Tikhonov regularization, total-variation
Cluitmans et al. (2018) [79] 3 canines BSP 192 ATM, EpiP Wavelet regularization
Zhou et al. (2018) [65] 7 VT patients ECG; BSP 120 OVA multi-linear regression; L1-norm regularization
Kalinin et al. (2019) [177] in silico BSP N/A EpiP, EndoP Tikhonov regularization
Kara et al. (2019) [60] in silico BSP 32/64/128/512/ 1024 Myo ischemia, EpiP Tikhonov regularization
Bear et al. (2020) [61] 4 canines/swines BSP 192/128 EpiP, ATM Tikhonov regularization
Rababah et al. (2021) [62] 6 canines/swines BSP 192/128 EpiP, ATM Tikhonov regularization
Schuler et al. (2021) [50] in silico BSP 200 EpiP, TMP, ATM Tikhonov regularization
Messnarz et al. (2004) [70] in silico BSP 62 TMP spatiotemporal regularization
@@
Erem et al. (2013) [76] 3 subjects BSP 120 EndoP, ATM, pacing sites temporal spline model
Erem et al. (2014) [77] in silico; 1 subject BSP 65 TMP, ATM temporal spline model
Zhou et al. (2016) [73] 7 AF patients ECG 208 atrial EndoP spatiotemporal frequency based model
Yao et al. (2017) [178] 4 MI patients BSP N/A MI, EGM spatiotemporal regularization
Coll et al. (2017) [74] in silico; 1 canine BSP 300, 354 T-wave alternans, EpiP, EndoP spatiotemporal spectral and frequency based model
Yang et al. (2018) [49] in silico; 4 swines BSP 208, 128 ATM spatial gradient sparse in frequency domain
Li et al. (2001) [93] in silico BSP 200 OVA model optimization based method
Tilg et al. (2002) [179] 2 patient ECG 55, 49 ATM model optimization based method
Calderero et al. (2005) [88] 1 canines BSP 711 activation wavefront level set
Liu et al. (2006) [48] in silico BSP 64/96/128/155/ 200 ATM biophysical model-based method
Han et al. (2008) [95] 4 rabbits BSP 53 ATM biophysical model-based method
Chavez et al. (2015) [89] 3 canines BSP N/A Myo ischemia iterative level set
Cluitmans et al. (2017) [85] 3 canines BSP 184-216 EpiP physiology-based regularization
Yang et al. (2017) [103] in silico; 4 PVC patients ECG 3 origin of PVC data-driven neural network
@@
Karoui et al. (2019) [110] in silico BSP N/A EpiP spatial adaptative time delay neural network
Alawad et al. (2019) [180] in silico; 3 patients ECG N/A OVA transfer learning
Gyawali et al. (2019) [181] 39 VT patients ECG N/A OVA sequential autoencoder
Cheng et al. (2021) [108] in silico BSP N/A TMP, MI shrinkage-thresholding network
Mu et al. (2021) [105] in silico BSP N/A TMP, MI, ATM shrinkage-thresholding + graph CNN
Camara (2021) et al. [107] in silico BSP 64 AF driver data-driven neural network
Karoui et al. (2021) [111] in silico BSP N/A atrial ATM data-driven neural network
Meister et al. (2021) [104] 16 swines ECG N/A ATM graph convolutional regression
Chen et al. (2022) [109] 5 swines BSP 239 EpiP, ATM, pacing sites data-driven neural network
Tenderini et al. (2022) [106] in silico ECG 9 EpiP PDE-aware deep learning model
@@
Jiang et al. (2024) [121] in silico; 3 patients BSP N/A EpiP, EndoP neural state-space modeling
Lian et al. (2024) [182] in silico; 4 patients ECG 9 MI frequency-enhanced network
Jiang et al. (2022) [83] 1 canine BSP 192 EpiP physics embedded graph CNN
Xie et al. (2023) [117] in silico BSP N/A EpiP physics-constrained network
Probabilistic estimation Serinagaoglu et al. (2006) [128] 1 canine BSP 771 EpiP Bayesian MAP
Wang et al. (2010) [183] in silico BSP N/A TMP, MI Bayesian MAP
Zhou et al. (2018) [127] in silico; 3 patients BSP 120 EndoP, pacing sites sparse Bayesian MAP
Zhou et al. (2018) [126] in silico; 3 VT patients BSP 120 EpiP spatio-temporal Bayesian MAP
Wang et al. (2009) [92] in silico; 1 MI patient BSP 700/330, 123 TMP physiological-constrained Kalman model
Huang et al. (2022) [132] in silico BSP N/A TMP data-driven Kalman filtering network
Wang et al. (2010) [184] 1 canine BSP N/A EpiP variational-form-based regularizers
Wang et al. (2011) [185] 1 canine BSP N/A EpiP variational regularizers
Xu et al. (2014) [186] in silico; 6 patients BSP N/A MI variational Bayesian
Figuera et al. (2016) [53] in silico BSP N/A atrial EpiP Bayesian MAP, etc
Ghimire et al. (2019) [142] in silico; 2 patients ECG N/A TMP Bayesian model
Ozkoc et al. (2021) [187] 1 canine BSP 192 EpiP Bayesian MAP
Li et al. (2024) [188] in silico ECG 10 MI multi-modal VAE
Table 4: Summary of the selected ECG inverse inference works for EP parameter estimation. HF: heart failure; APD: action potential duration; MS: Mitchell-Schaeffer; AP: Aliev-Panfilov.
Source & Publish year Data Signal Target EP model Inverse inference method
Deterministic estimation Chinchapatnam et al. (2008) [100] in silico; 1 patient N/A local CV; conductivity Eikonal adaptive zonal iterative model
Sermesant et al. (2012) [101] 2 patients ECG local CV, EAS Eikonal model-based regularization
Potse et al. (2014) [189] 2 HF patients ECG EAS, conductivity, etc Ten Tusscher and Panfilov model-based regularization
Kahlmann et al. (2017) [190] in silico ECG Purkinje network Eikonal parallel simplex optimization
Sanchez et al. (2018) [191] 6 patients ECG EAS, conductivity, etc Ten Tusscher and Panfilov iterative method
Lee et al. (2019) [192] in silico ECG global CV reaction-Eikonal rule based model
Pezzuto et al. (2021) [193] 11 HF patients ECG global CV, EAS Eikonal propagation model
Grandits et al. (2021) [98] in silico; 1 patient ECG EAS Eikonal topological gradient
Costa et al. (2022) [8] 6 pigs ECG global CV, APD, conductivity reaction-Eikonal total activation time fit
Grandits et al. (2023) [99] 1 patient ECG local CV, EAS, ATM Eikonal Geodesic backpropagation
Neumann et al. (2016) [114] in silico ECG global CV, Purkinje network Eikonal reinforcement learning
Giffard et al. (2018) [115] 20 patients BSP local CV, pacing site, ATM MS transfer learning
Herrero et al. (2022) [119] in silico; in vitro ECG APD, excitability and diffusion coefficients AP PINN
Ye et al. (2023) [118] in silico ECG tissue excitability AP PINN
@@
Probabilistic estimation Zettinig et al. (2014) [46] in silico ECG electrical diffusivity MS statistical polynomial regression
Gillette et al. (2021) [2] 1 patient ECG global CV, Purkinje network, APD, conductivity reaction-Eikonal Saltelli sampling
Camps et al. (2021) [1] in silico ECG global CV, EAS Eikonal Monte Carlo Bayesian inf